OpenStudy (anonymous):

The length of a side of a square is (4x − 1) yd. Find the area of the square in terms of the variable x. _____________yd2

jimthompson5910 (jim_thompson5910):

A = s^2 A = (4x-1)^2 A = (4x-1)(4x-1) A = 4x*4x-4x*1-1*4x-1*(-1) A = 16x^2-4x-4x+1 A = 16x^2-8x+1 So the area is 16x^2-8x+1 square yards